# Activation of anomalous Hall effect and orbital magnetization by domain walls in altermagnets

### v1

@@ Abstract
Altermagnets are an emerging class of unconventional antiferromagnets, characterized by a Néel ordering that does not break the translation symmetry of the underlying lattice. Depending on the orientation of the Néel vector, the anomalous Hall effect (AHE) may or may not exist. In the so-called pure altermagnets, AHE is forbidden by the magnetic symmetry. Here, we demonstrate that in pure altermagnets, the domain walls can lift the symmetry constraints, thereby activating the AHE and orbital magnetization. Taking a representative example of a rutile-lattice tight-binding minimal model in slab geometry, we use the linear response theory to demonstrate the emergence of the domain wall AHE, finding that it is closely related with the orbital magnetization, while the spin magnetization does not play a significant role. Using Landau theory, we argue that while for a random arrangement of $\pi$ domain walls, the contributions from the individual domain walls will cancel one another, an external magnetic field will favor domain-wall arrangements with specific chirality giving rise to a net AHE signal. Using group theory, we discuss how these findings can be generalized straightforwardly to certain other classes of altermagnets. Our work reveals a crucial role of the domain walls in the understanding of the Hall transport and orbital magnetism of altermagnets.
@@ I Introduction
Figure 1: Geometrical setup of a film periodic in the $yz$-plane featuring an open boundary condition along the $x$-axis. The double-sided arrows denote local Néel vectors that form a Bloch-type domain wall along $x$.
@@ II Model
Figure 4: Bulk band structures for (a) $\vec{N}_{i}=\hat{z}$ and (b) $\vec{N}_{i}=\hat{y}$. A notable distinction between the two panels is the band anticrossings along the AZ line, which is plotted in a different color. In (a), the band crossings along AZ are protected by $z$-mirror and are a part of a Weyl nodal loop in the $k_{z}=\pi$ plane. In (b), the Néel vector lowers the symmetry, leading to the gap opening. The symmetry protection of the Weyl nodal lines such as those visible along $\Gamma$Z and $MA$, is discussed in the main text. The model parameters used are $t_{1}=-0.2,t_{2}=0.025,t^{\prime}_{2}=-0.35,t_{d}=0.15$, $J=0.2,\lambda=-0.025$ and $\lambda^{\prime}=0.0125,$ see main text.
Figures 4(a) and (b) show the band structure of the model for $\vec{N}=\hat{z}$ and $\vec{N}=\hat{y}$, respectively. As visible in (a), the band structure supports Weyl nodal lines, such as those along MA and $\Gamma$Z. On the $k_{z}=\pi$ plane, there are also Weyl nodal loops which intersect along AZ and manifest as two band-crossing points. The symmetry protection of these Weyl nodal structures and their physical implications, e.g. a rise of unconventional antichiral surface states, have been discussed in Ref. [17, 18]. Particularly relevant for our AHE considerations is the Weyl nodal loops in the $k_{z}=\pi$ plane, which are protected by mirror $\mathcal{M}_{z}$ symmetry [17]. When $\vec{N}=\hat{y}$, the mirror protection is removed, and the band-crossing points are gapped out, as can be seen in Fig. 4(b). These anticrossings support a large concentration of local Berry curvature which results in a nontrivial dependence of the AHE on the chemical potential, as will be seen in Sec. II.1. Compared with AZ, the Weyl nodal lines along $\Gamma$Z and MA seem to be intact upon the change in the direction of the Néel vector. A more detailed analysis reveals that, indeed, the Weyl nodal line along MA is protected by a combination of a nonsymmorphic symmetry and an antiunitary symmetry (see Appendix A.) However, the nodal line along $\Gamma$Z is not protected by any symmetry and can be gapped out by a symmetry-conforming perturbation (see Appendix A.)
@@ II.1 Anomalous Hall effect and orbital magnetization in collinear bulk
In this Section, we will consider the AHE in the bulk with a fixed direction of the Néel vector, i.e. when the DWs are absent, for two cases: (i) $\vec{N}=\hat{z}$ and (ii) for $\vec{N}=\hat{y}$. Case (i) corresponds to a pure altermagnet, where AHE is forbidden by two mirror symmetries: $S_{1}=\{\mathcal{M}_{x}|\frac{1}{2}\frac{1}{2}\frac{1}{2}\}$ and $S_{2}=\{\mathcal{M}_{z}|000\}$. $S_{1}$ requires the time-reversal-odd Hall pseudovector $\vec{\sigma}_{H}=(\sigma^{H}_{yz},\sigma^{H}_{zx},\sigma^{H}_{xy})^{T}$ to be in the $x$-direction, where $\sigma_{ij}^{H}$ is the Hall conductivity in the $ij-$plane. Meanwhile, $S_{2}$ requires $\vec{\sigma}_{H}$ to be in the $z$-direction. As a result, $\vec{\sigma}_{H}=0$.
@@
where we have assumed the spin $g$-factor to be 2 and the unit such that the bare electron mass $m_{e}$ is 1. $(v_{j})^{nm}_{\vec{k}}=\bra{u_{\vec{k}n}}v_{j}\ket{u_{\vec{k}m}}$, |$\ket{u_{\vec{k}m}}$ is the Bloch eigenvector, and $v_{j}$ is the $j$’th Cartesian component of the velocity operator, given by $\frac{1}{\hbar}\frac{\partial\mathcal{H}}{\partial k_{i}}$. For the choice of the hopping integrals, the effective mass of electrons is comparable with the bare mass, which allows a direct comparison between the numerical results for $M_{\text{orb},i}$ and $M_{\text{sp},i}$. [^2] The effective mass of the electrons is of the order of $\hbar^{2}/a^{2}W$, where $W$ denotes the bandwidth, and $a$ is the lattice constant. This means that one can vary the tight-binding parameters in order to change $W$ and the effective mass. Assuming the unit where $\hbar=a=m_{e}=1$, the bandwidth $W$ as seen in Fig.4 produces the effective mass comparable with the bare mass $m_{e}$. Therefore, we can indeed compare the order of magnitude between $\vec{M}_{\rm orb}$ and $\vec{M}_{\rm sp}$ with the choice of the tight-binding parameters.
@@ III Anomalous Hall effect from altermagnetic domain walls
To study the AHE induced by altermagnetic DWs, we consider a Bloch DW of the Néel vector in the slab geometry as shown in Fig. 1. We consider the DW profile where the Néel vector in the $i$-th unit cell is given by $\vec{N}_{i}$, for $-L_{x}/2<i<L_{x}/2$, and $L_{x}$ is the thickness of the slab. We use the following Ansatz for the Néel vector profile: $\vec{N}_{i}=(0,\sin\theta_{i},\cos\theta_{i})^{T}$. For the left-hand-side domain, $i\leq-w_{\rm dw}/2$, $\theta_{i}=0$, and for the right-hand-side domain $i\geq w_{\rm dw}/2$, $\theta_{i}=\pi$. In the DW region, $-w_{\rm dw}/2<i<w_{\rm dw}/2$, $\theta_{i}=\pi(2i+w_{\rm dw})/2w_{\rm dw}$, where $w_{\rm dw}$ characterizes the width of the DW. We then couple the conduction electrons hopping on the rutile-lattice slab geometry with this DW profile. Reference [18] studies the electronic spectrum in the presence of such DWs, uncovering the presence of unconventional bound states at the DWs. Here, we focus on how the DW gives rise to a nontrivial anomalous Hall response.
@@ III.1 Anomalous Hall effect from altermagentic domain walls
We compute the DW Hall conductivity $\sigma^{H,DW}_{ij}$ in the presence of the DW using Kubo formula and the eigenvectors of the Hamiltonian matrix for the slab-geometry problem. The full expression for the Hall conductivity can be found in the Appendix B (see also Ref. [37]). We will only show the results for $\sigma^{H,DW}_{yz}$ and $\sigma^{H,DW}_{xy}$, since $\sigma^{H,DW}_{zx}$ is forbidden by the symmetry operation of a mirror $\mathcal{M}_{y}$ followed by the time reversal and a fractional translation along the $y$-direction.
@@
The result for $\sigma^{H,avg}_{yz}$ is shown as the dashed line in Fig. 6(b). We find that the site averaging works remarkably well as compared to the full calculation. Based on this, one can effectively view the sign-changing behavior of $\sigma_{yz}^{H,DW}$ near $\mu=-0.75$ as the result of band anticrossings akin to the collinear-bulk cases, but here the anticrossings are caused by the DW texture. In general, we expect the site averaging to work better for a smoother DW profile since the DW-induced potential felt by the electrons varies more slowly in space. A notable consequence is that, in a smoother and correspondingly wider DW, there are more unit cells that contribute a nonzero $\sigma^{H}_{ab}(\vec{N}_{i})$ in the right-hand side of Eq. (7). In other words, for a fixed $L_{x}$, a wider DW with a larger $w_{\rm dw}$ has nonzero contributions from more $i$-th sites in Eq.(7). Consequently, the Hall conductivity is larger. Therefore, DWs with a larger width are expected to induce a larger DW AHE. Next, we study how the DW AHE is related to the orbital and the spin magnetization.
@@ III.2 Relation to the orbital and spin magnetization
$M_{\text{sp},i}$ $=$ $-\frac{e\hbar}{2}\frac{1}{L_{x}+1}\sum_{n}\int\frac{d^{2}q}{(2\pi)^{2}}f(E_{\vec{q}n})\bra{u_{\vec{q}n}}\sigma_{i}\ket{u_{\vec{q}n}},$ (8)
@@ III.4 AHE from mutli-domain-wall profiles
where $\sigma_{yz}^{H,DW}$ is the Hall conductivity from the slab-geometry calculation corresponding to the slab thickness $L_{x}$ and a DW width $w_{\rm dw}$. This is expected to be a good approximation, assuming that the Hall contributions from the constituent DWs do not strongly interfere with one another, and thus these contributions simply add up. We note that the final outcome depends on $\rho_{\rm dw}$ and $w_{\rm dw}$. The latter is intrinsic to each DW, which determines the product $\sigma_{yz}^{H,DW}L_{x}$. This product is expected to converge for a sufficiently large $L_{x}$. In other words, we expect the AHE contribution from each DW to be mainly determined by the properties of the DW itself, which is the width $w_{\rm dw}$ in this case. For $L_{x}=100a$ and $w_{\rm dw}=20a$, $\sigma_{yz}^{H,DW}$ is of the order of $10^{-1}e^{2}/ha$; see Fig. 6(b). For $\rho_{\rm dw}=1/1000\,a$, i.e. one DW per 1000 lattice constant, the total Hall conductivity is of the order of $10^{-2}e^{2}/ha$. For $a=1$ nm, the final value of the Hall conductivity is of the order of magnitude between $1$ S/cm and $10$ S/cm, which is an appreciable value. In the next section, we return to the demonstration of how an applied magnetic field may indeed select the DW patterns that lead to a nonzero $\sigma_{yz}^{H,DW}$.
Figure 8: The four possible DW-pair combinations that are expected to arise in multiple-parallel-DW profiles. They differ by the combination of the sign of $N_{y}$ at the two DWs. (a) $N_{y}$ component at the two DWs has (+, +) sign. (b) (+, --) sign. (c) (--, +) sign. (d) (--, --) sign. If an applied magnetic field favors a positive $N_{y}$, then the combination (a) has the lowest energy.
@@ IV Insights from Landau theory
$mmm$ $B_{1g}$ $\alpha_{1}N_{x}M_{y}+\alpha_{2}N_{y}M_{x}$ along $\hat{z}$
@@ V Generalization beyond rutile altermagnets
We find three crystallographic point groups in three dimensions which support pure altermagnetism and the bilinear couplings simultaneously; see Table 1. This table is adapted from a recent work on Landau theories for altermagents [45], and we will describe how to use it in the following. The first column is the crystallographic point group of the underlying lattice. To describe the physical meaning of the second column, we first recall that an altermagnetic ordering does not break the translational symmetry of the lattice. It is instructive to picture how the staggering dipole moments associated with the Néel order occupy and divide the sublattices in each unit cell into two groups. We define a sign-alternating pattern by assigning the value +1 to one group and --1 to the other group. It is essential that a translation or an inversion operation does not interchange the sublattices between the two groups to avoid the Kramers’ degeneracy. Otherwise, we will get the conventional antiferromagnetism, i.e. no momentum-dependent band splitting. Nevertheless, the two groups may be interchanged by other operations, e.g. a rotation operation of the crystallographic point group  [^4] In practice, one needs to consider symmetry operations of the space group, but each of such operations can be identified unambiguously with an operation of the point group.. As a result, such a rotation reverses the sign of the sign-alternating pattern. While some operations of the point group change the sign of the sign-alternating pattern, other operations do not. This means that the sign-alternating pattern transforms according to a one-dimensional irreducible representation of the point group which defines $\Gamma_{S}$. $\Gamma_{S}$ varies, depending on the type of Wykoff positions taken by the sublattices. For rutiles, the point group is $4/mmm$. Within each unit cell, the A sublattice constitutes the first group, while the B sublattice forms the second group. A four-fold rotation is an example that interchanges the two groups. The sign-alternating pattern transforms according to $\Gamma_{S}=B_{2g}$. Accordingly, the bilinear coupling can be read off from the third column of the table and is indeed given by Eq. (14). Finally, the fourth column specifies the high-symmetry direction for $\vec{N}$ corresponding to the pure altermagnetism. Table 1 demonstrates the presence of multiple examples of pure altermagnets where the bilinear couplings are possible. Our results for the rutile altermagnets can then be straightforwardly generalized for these systems.
@@ VI Summary
Our work demonstrates that in pure altermagnets the symmetry constraint, which forces the AHE to vanish, is lifted by magnetic DWs, thereby activating a DW AHE and the orbital magnetization. In the presence of many parallel DWs, while a random arrangement of the DW chirality leads to a trivial DW AHE, we argue using Landau theory how an external magnetic field selects a chirality population that allows the AHE contribution from each DW to add up instead of canceling each other off. Our work also uncovers a crucial role of the orbital magnetization, in that it is comparable to or, in some cases, even larger than the spin magnetization. In addition, its orientation correctly indicates the Hall plane where the DW AHE is the greatest, whereas the spin magnetization fails to do so. Our work calls for experimental investigations to confirm such an important role of altermagetic DWs in Hall transport and orbital magnetism phenomena. Future studies are also needed to firmly establish the dominant effect of the orbital magnetism uncovered here in altermagnets. This last point is further reinforced by a similar observation of a dominant orbital magnetism in MnTe in a recent first-principle study [46].

### v2

@@ Abstract
Altermagnets are an emerging class of unconventional antiferromagnets, characterized by a Néel ordering that does not break the translation symmetry of the underlying lattice. Depending on the orientation of the Néel vector, the anomalous Hall effect (AHE) may or may not exist. In the so-called pure altermagnets, AHE is forbidden by the magnetic symmetry. Here, we demonstrate that in pure altermagnets, the domain walls can lift the symmetry constraints, thereby activating the AHE and orbital magnetization. Taking a representative example of a rutile-lattice tight-binding minimal model in slab geometry, we use the linear response theory to demonstrate the emergence of the domain wall AHE, finding that it is closely related with the orbital magnetization, while the spin magnetization does not play a significant role. Using Landau theory, we argue that while for a random arrangement of $\pi$ domain walls, the contributions from the individual domain walls will cancel one another, an external magnetic field will favor domain-wall arrangements with specific chirality giving rise to a net AHE signal. Using group theory, we discuss how these findings can be generalized straightforwardly to certain other classes of altermagnets. Our work reveals a crucial role of the domain walls in the understanding of the Hall transport and orbital magnetism of altermagnets. Our work generally calls for a rigorous analysis of Hall-transport data for an altermagnet, ideally in conjunction with an imaging data, in order to unambiguously assign an observed Hall effect to the magnetic domains or to the domain walls.
@@ I Introduction
Figure 1: Geometrical setup of a film periodic in the yzyz-plane featuring an open boundary condition along the $x$-axis. The double-sided arrows denote local Néel vectors that form a Bloch-type domain wall along $x$.
@@ II Model
[…]
@@ II.1 Anomalous Hall effect and orbital magnetization in collinear bulk
In this Section, we will consider the AHE in the bulk with a fixed direction of the Néel vector, i.e. when the DWs are absent, for two cases: (i) $\vec{N}=\hat{z}$ and (ii) for $\vec{N}=\hat{y}$. Case (i) corresponds to a pure altermagnet, where AHE is forbidden by two mirror symmetries: $S_{1}=\{\mathcal{M}_{x}|\frac{1}{2}\frac{1}{2}\frac{1}{2}\}$ and $S_{2}=\{\mathcal{M}_{z}|000\}$. $S_{1}$ requires the time-reversal-odd Hall pseudovector $\vec{\sigma}_{H}=(\sigma^{H}_{yz},\sigma^{H}_{zx},\sigma^{H}_{xy})^{T}$ to be in the $x$-direction, where $\sigma_{ij}^{H}$ is the Hall conductivity in the ijij-plane. Meanwhile, $S_{2}$ requires $\vec{\sigma}_{H}$ to be in the $z$-direction. As a result, $\vec{\sigma}_{H}=0$.
@@
[…]
@@ III Anomalous Hall effect from altermagnetic domain walls
To study the AHE induced by altermagnetic DWs, we consider a Bloch DW of the Néel vector in the slab geometry as shown in Fig. 1. We consider the DW profile where the Néel vector in the $i$-th unit cell is given by $\vec{N}_{i}$, for $L_{x}/2<i<L_{x}/2$, and $L_{x}$ is the thickness of the slab. We use the following Ansatz for the Néel vector profile: $\vec{N}_{i}=(0,\sin\theta_{i},\cos\theta_{i})^{T}$. For the left-hand-side domain, $i\leq-w_{\rm dw}/2$, $\theta_{i}=0$, and for the right-hand-side domain $i\geq w_{\rm dw}/2$, $\theta_{i}=\pi$. In the DW region, $w_{\rm dw}/2<i<w_{\rm dw}/2$, $\theta_{i}=\pi(2i+w_{\rm dw})/2w_{\rm dw}$, where $w_{\rm dw}$ characterizes the width of the DW. We then couple the conduction electrons hopping on the rutile-lattice slab geometry with this DW profile. Reference [18] studies the electronic spectrum in the presence of such DWs, uncovering the presence of unconventional bound states at the DWs. Here, we focus on how the DW gives rise to a nontrivial anomalous Hall response.
@@ III.1 Anomalous Hall effect from altermagentic domain walls
We compute the DW Hall conductivity $\sigma^{H,DW}_{ij}$ in the presence of the DW using Kubo formula and the eigenvectors of the Hamiltonian matrix for the slab-geometry problem. The full expression for the Hall conductivity can be found in the Appendix B (see also Ref. [37]). We will only show the results for $\sigma^{H,DW}_{yz}$ and $\sigma^{H,DW}_{xy}$, since $\sigma^{H,DW}_{zx}$ is forbidden by the symmetry operation of two-fold rotation $C_{2y}$ followed by the time reversal and a fractional translation along the $y$-direction.
@@
The result for $\sigma^{H,avg}_{yz}$ is shown as the dashed line in Fig. 6(b). We find that the site averaging works remarkably well as compared to the full calculation. Based on this, one can effectively view the sign-changing behavior of $\sigma_{yz}^{H,DW}$ near $\mu=-0.75$ as the result of band anticrossings akin to the collinear-bulk cases, but here the anticrossings are caused by the DW texture. In general, we expect the site averaging to work better for a smoother DW profile since the DW-induced potential felt by the electrons varies more slowly in space. A notable consequence is that, in a smoother and correspondingly wider DW, there are more unit cells that contribute a nonzero $\sigma^{H}_{ab}(\vec{N}_{i})$ in the right-hand side of Eq. (7). In other words, for a fixed $L_{x}$, a wider DW with a larger $w_{\rm dw}$ has nonzero contributions from more $i$-th sites in Eq.(7). Consequently, the Hall conductivity is larger. Therefore, DWs with a larger width are expected to induce a larger DW AHE. We note also that the agreement between $\sigma^{H,avg}_{yz}$ and $\sigma^{H,DW}_{yz}$ provides strong evidence that the Hall effect is indeed activated by the DW, while bulk regions far from the DW do not contribute to the Hall effect, as expected from the altermagnetic symmetry of each domain. Through the definition of $\sigma^{H,avg}_{yz}$, it is also clear that an experimental setup where the Hall voltage contacts cover the whole vertical x-direction of the slab is needed in order to observe the slab-averaged DW Hall effect given by $\sigma^{H,avg}_{yz}$ and equivalently the $\sigma^{H,DW}_{yz}$. Next, we study how the DW AHE is related to the orbital and the spin magnetization.
@@ III.2 Relation to the orbital and spin magnetization
[…]
@@ III.4 AHE from mutli-domain-wall profiles
where $\sigma_{yz}^{H,DW}$ is the Hall conductivity from the slab-geometry calculation corresponding to the slab thickness $L_{x}$ and a DW width $w_{\rm dw}$. This is expected to be a good approximation, assuming that the Hall contributions from the constituent DWs do not strongly interfere with one another, and thus these contributions simply add up. We note that the final outcome depends on $\rho_{\rm dw}$ and $w_{\rm dw}$. The latter is intrinsic to each DW, which determines the product $\sigma_{yz}^{H,DW}L_{x}$. This product is expected to converge for a sufficiently large $L_{x}$. In other words, we expect the AHE contribution from each DW to be mainly determined by the properties of the DW itself, which is the width $w_{\rm dw}$ in this case. For $L_{x}=100a$ and $w_{\rm dw}=20a$, $\sigma_{yz}^{H,DW}$ is of the order of $10^{-1}e^{2}/ha$; see Fig. 6(b). For $\rho_{\rm dw}=1/1000\,a$, i.e. one DW per 1000 lattice constant, the total Hall conductivity is of the order of $10^{-2}e^{2}/ha$. For $a=1$ nm, the final value of the Hall conductivity is of the order of magnitude between $1$ S/cm and $10$ S/cm, which is an appreciable value. Our work reveals a non-negligible role of DWs in understanding Hall transport phenomena in altermagnets, which is particularly crucial for pure altermagnets since the bulk magnetic domains do not support AHE. Therefore, our work generally calls for a proper analysis of the Hall transport data, ideally simultaneously with magnetic imaging data, in order to correctly attribute the observed Hall effect to the bulk magnetic domains or to DWs. In the next section, we return to the demonstration of how an applied magnetic field may indeed select the DW patterns that lead to a nonzero $\sigma_{yz}^{H,DW}$.
Figure 8: The four possible DW-pair combinations that are expected to arise in multiple-parallel-DW profiles. They differ by the combination of the sign of $N_{y}$ at the two DWs. (a) $N_{y}$ component at the two DWs has (+, +) sign. (b) (+, -) sign. (c) (-, +) sign. (d) (-, -) sign. If an applied magnetic field favors a positive $N_{y}$, then the combination (a) has the lowest energy.
@@ IV Insights from Landau theory
mmmmmm $B_{1g}$ $\alpha_{1}N_{x}M_{y}+\alpha_{2}N_{y}M_{x}$ along $\hat{z}$
@@ V Generalization beyond rutile altermagnets
We find three crystallographic point groups in three dimensions which support pure altermagnetism and the bilinear couplings simultaneously; see Table 1. This table is adapted from a recent work on Landau theories for altermagents [45], and we will describe how to use it in the following. The first column is the crystallographic point group of the underlying lattice. To describe the physical meaning of the second column, we first recall that an altermagnetic ordering does not break the translational symmetry of the lattice. It is instructive to picture how the staggering dipole moments associated with the Néel order occupy and divide the sublattices in each unit cell into two groups. We define a sign-alternating pattern by assigning the value +1 to one group and -1 to the other group. It is essential that a translation or an inversion operation does not interchange the sublattices between the two groups to avoid the Kramers’ degeneracy. Otherwise, we will get the conventional antiferromagnetism, i.e. no momentum-dependent band splitting. Nevertheless, the two groups may be interchanged by other operations, e.g. a rotation operation of the crystallographic point group  [^4] In practice, one needs to consider symmetry operations of the space group, but each of such operations can be identified unambiguously with an operation of the point group.. As a result, such a rotation reverses the sign of the sign-alternating pattern. While some operations of the point group change the sign of the sign-alternating pattern, other operations do not. This means that the sign-alternating pattern transforms according to a one-dimensional irreducible representation of the point group which defines $\Gamma_{S}$. $\Gamma_{S}$ varies, depending on the type of Wykoff positions taken by the sublattices. For rutiles, the point group is $4/mmm$. Within each unit cell, the A sublattice constitutes the first group, while the B sublattice forms the second group. A four-fold rotation is an example that interchanges the two groups. The sign-alternating pattern transforms according to $\Gamma_{S}=B_{2g}$. Accordingly, the bilinear coupling can be read off from the third column of the table and is indeed given by Eq. (14). Finally, the fourth column specifies the high-symmetry direction for $\vec{N}$ corresponding to the pure altermagnetism. Table 1 demonstrates the presence of multiple examples of pure altermagnets where the bilinear couplings are possible. Our results for the rutile altermagnets can then be straightforwardly generalized for these systems.
@@ VI Summary
Our work demonstrates that in pure altermagnets the symmetry constraint, which forces the AHE to vanish, is lifted by magnetic DWs, thereby activating a DW AHE and the orbital magnetization. In the presence of many parallel DWs, while a random arrangement of the DW chirality leads to a trivial DW AHE, we argue using Landau theory how an external magnetic field selects a chirality population that allows the AHE contribution from each DW to add up instead of canceling each other off. Our work also uncovers a crucial role of the orbital magnetization, in that it is comparable to or, in some cases, even larger than the spin magnetization. In addition, its orientation correctly indicates the Hall plane where the DW AHE is the greatest, whereas the spin magnetization fails to do so. Our work calls for experimental investigations to confirm such an important role of altermagetic DWs in Hall transport and orbital magnetism phenomena. Due to the important role of DWs in pure altermagnets demonstrated here, future Hall-transport studies should undertake a rigorous analysis of the Hall data, ideally in conjunction with the information on the precise magnetic structure, in order to unambiguously assign an observed Hall effect to the bulk magnetic domains or to the DWs. Our work also motivates future studies of DW effects in mixed altermagnets, where bulk magnetic domains are already AHE-active and are accompanied by a weak ferromagnetism. It is probable that, in a field-sweep Hall measurement, the Hall contribution from bulk magnetic domains cancel each other off near the coercive magnetic field, exposing the DW contribution—a scenario proposed in the ferromagnet CeAlSi [38]. Future studies are also needed to firmly establish the dominant effect of the orbital magnetism uncovered here in altermagnets. This last point is further reinforced by a similar observation of a dominant orbital magnetism in MnTe in a recent first-principle study [46].